# Schroedinger’s Threshold: When the AUC doesn’t predict Accuracy

### v1

@@ 1.   Introcuction
Yet, an issue is that AUC yields a academic view on model power. In the real-world, we cannot forgo model calibration, as we ultimately have to make decisions. In our example of text faithfulness, there are clear ramifications of different decision thresholds: with a false-positive we run a risk of releasing false or even harmful output; a false-negative may lead to censoring of good system output.
In this paper, we show that such important real world considerations tend to be neglected by the AUC, and find that its theoretical perspective on system performance may not align with actual performance in applications. Our findings indicate that a main factor for this lies in the diversity of model score and data distributions. Indeed, we argue that AUC should not be used as a sole measure for model evaluation and benchmarking, particularly when models and data are diverse.
@@ 3.   Experimental setup
Models are also diverse. Some are NLI-based (a closely related task), while others employ elaborate scoring techniques, e.g., by analyzing a cross-product of sentences. As in TRUE, we employ T5 ANLI (Honovich et al., 2022) which is a T5-11B (Raffel et al., 2020) LLM trained on ANLI Nie et al. (2020). SummacZS (Laban et al., 2022) evaluates an NLI model on sentence pairs and averages maximum entailment probabilities, and Q2 (Honovich et al., 2021) integrates a question-answering step.
@@ 3.1.   Measurement of expected accuracy
Table 1: AUC evaluation (x100).
@@
Table 2: Expected accuracy evaluation (x100).
@@ 4.3.   Studying score distribution
We create two groups of models: those that obtain a better rank according to accuracy, and those that obtain a worse rank. The histograms are shown in Figure 4. We see that models that are relatively more negatively affected by calibration tend to show more skewed distributions. The scores of the better models, however, seem more balanced and also exhibit a smaller average variance (Table 3: avg. variance of better metrics=0.05; avg. variance of worse metrics=0.12).
@@
Table 4: Different modes of calibration, varying calibration method and training data. Mean performance over all data sets. In each group of three lines: The first line is calibration via logistic regression, the second line is isotonic regression, and the third is decision stump. Note the assumptions on data availibility: Indata requires annotated in-domain in-task training; InDomain needs in-domain training; XDomain lessens this dependence, and OutDomain is the most general setup. ***: A data subset (PAWS) is calibrated InData (instead of Indomain), since it is the only data set of domain paraphrase/wiki.
@@ 5.1.   Effect of calibration technique
We denote the cross-domain setup from the section before as Xdomain. Additionally, we introduce the arguably hard setup of OutDomain which is interesting since it only allows training on out-domain training data and thus tests the transfer to new domains. Other setups are InDomain that allows calibration only in in-domain training data, and InData, where we have in-domain and in-task training data, where we would naturally expect the best although much less generalizable performance. For InData, we estimate performance on a random 80/20 train/test split of a data set, averaged over 100 repetitions.
@@
The intuitive logistic curve calibration is by far not the only possible calibration method one. In fact, it has also been criticized Silva Filho et al. (2023), e.g., due to observed over-confidence effects.

### v2

@@ 1.   Introcuction
Yet, an issue is that AUC has an academic view on model power. In a “real-world” application, we cannot forgo model calibration, as we ultimately have to make decisions. In our example of text faithfulness, there are clear ramifications of different decision thresholds: with a false-positive we run a risk of releasing false or even harmful output; a false-negative may lead to censoring of good system output.
In this paper, we show that such important real world considerations tend to be neglected by the AUC, and find that its theoretical perspective on system performance may not align with actual performance in applications. Our findings indicate that a main factor for this lies in the diversity of model score and data distributions. We thus argue that AUC should not be used as a sole measure for model evaluation and benchmarking, particularly when models and data are diverse.
@@ 3.   Experimental setup
Models are also diverse. Some are NLI-based (a closely related task), while others employ elaborate scoring techniques, e.g., by analyzing a cross-product of sentences. As in TRUE, we employ ANLI (Honovich et al., 2022) which is a T5-11B (Raffel et al., 2020) LLM trained on ANLI Nie et al. (2020). SummacZS (Laban et al., 2022) evaluates an NLI model on sentence pairs and averages maximum entailment probabilities, and Q2 (Honovich et al., 2021) integrates a question-answering step.
@@ 3.1.   Measurement of expected accuracy
Table 1: AUC evaluation (x100). Bold integers: rank of model on a data set.
@@
Table 2: Expected accuracy evaluation (x100). Bold integers: rank of model on a data set.
@@ 4.3.   Studying score distribution
We create two groups of models: those that obtain a better rank according to accuracy, and those with a worse rank. The histograms are shown in Figure 4. We see that models that are relatively more negatively affected by calibration tend to show more skewed distributions. The scores of the better models seem more balanced and exhibit a smaller average variance (Table 3: avg. variance of better metrics=0.05; avg. variance of worse metrics=0.12).
@@
Table 4: Different modes of calibration, varying calibration method and training data. Mean performance over all data sets. In each group of three lines: The first line is calibration via logistic regression, the second line is isotonic regression, and the third is decision stump. Note the assumptions on data availibility: Indata requires annotated in-domain in-distribution training; InDomain assumes availability of in-domain training data; XDomain assumes training data from different domains (including the testing domain), OutDomain assumes the testing data set is from a different domain than the training data, OutData assesses the utility of a data set to serve as calibration/training data set. ***: A data subset (PAWS) is calibrated InData (instead of Indomain), since it is the only data set of domain paraphrase/wiki.
@@ 5.1.   Effect of calibration technique
We denote the cross-domain setup from the section before as Xdomain. Additionally, we introduce the arguably harder setup of OutDomain that tests the transfer to new domains: e.g., if a testing data set is from domain ‘summarization’ we calibrate the model on ‘dialog’ and ‘paraphrase’ data sets. Other setups are InDomain that allows calibration only on in-domain data. For InData we estimate performance on a random 80/20 train/test split of a data set, averaged over 100 repetitions. Finally, in OutData only the data set itself is used to calibrate a model, and generalization is measured on all other data sets (OutData measures the utility of a data set to serve as a development set for calibration).
@@
The intuitive logistic curve calibration is by far not the only possible calibration method. In fact, it has also been criticized Silva Filho et al. (2023), e.g., due to observed over-confidence effects.